40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 349 of 350  Not logged in ELOG logo
ID Date Author Type Category Subject
  64   Mon Nov 5 22:24:38 2007 Andrey, SteveOmnistructureVACPumping down goes smoothly

We (Steve and Andrey) started pumping down at 3.25PM today. At 9 PM we turned off the rotary pump, and turned on turbomolecular pumps.

By 10.10PM we reached the pressure 1 milliTorr, and the current status is "Vacuum Normal". We leave the turbopumps on for the night, and as it is pretty late for Steve, we are going home.

P.S. Steve was very displeased with the standard selection of "Type" of messages, he would like to extend that list.
  63   Mon Nov 5 14:44:39 2007 waldmanUpdateOMCPZT response functions and De-whitening
The PZT has two control paths: a DC coupled path with gain of 20, range of 0 to 300 V, and a pair of 1:10 whitening filters, and an AC path capacitively coupled to the PZT via a 0.1 uF cap through a 2nd order, 2 kHz high pass filter. There are two monitors for the PZT, a DC monitor which sniffs the DC directly with a gain of 0.02 and one which sniffs the dither input with a gain of 10.

There are two plots included below. The first measures the transfer function of the AC monitor / AC drive. It shows the expected 2 kHz 2d order filter and an AC gain of 100 dB, which seems a bit high but may be because of a filter I am forgetting. The high frequency rolloff is the AA and AI filters kicking in which are 3rd order butters at 10 kHz.

The second plot is the DC path. The two traces show the transfer function of DC monitor / DC drive with and with an Anti-dewhitening filter engaged in the DC drive. I fit the antidewhite using a least squares routine in matlab constrained to match 2 poles, 2 zeros, and a delay to the measured complex filter response. The resulting filter is (1.21, 0.72) : (12.61, 8.67) and the delay was f_pi = 912 Hz. The delay is a bit lower than expected for the f_pi = 3 kHz delay of the AA, AI, decimate combination, but not totally unreasonable. Without the delay, the filter is (1.3, 0.7) : (8.2, 13.2) - basically the same - so I use the results of the fit with delay. As you can see, the response of the combined digital AntiDW, analog DW path is flat to +/- 0.3 dB and +/- 3 degrees of phase.

Note the -44 dB of DC mon / DC drive is because the DC mon is calibrated in PZT Volts so the TF is PZT Volts / DAC cts. To calculate this value: there are (20 DAC V / 65536 DAC cts)* ( 20 PZT V / 1 DAC V) = -44.2 dB. Perfect!

I measured the high frequency response of the loop DC monitor / DC drive to be flat.
Attachment 1: 07110_DithertoVmonAC_sweep2-0.png
07110_DithertoVmonAC_sweep2-0.png
Attachment 2: 071105_LSCtoVmonDC_sweep4-0.png
071105_LSCtoVmonDC_sweep4-0.png
Attachment 3: 07110_DithertoVmonAC_sweep2.pdf
07110_DithertoVmonAC_sweep2.pdf 07110_DithertoVmonAC_sweep2.pdf
Attachment 4: 071105_LSCtoVmonDC_sweep4.pdf
071105_LSCtoVmonDC_sweep4.pdf 071105_LSCtoVmonDC_sweep4.pdf
  62   Mon Nov 5 07:29:35 2007 ranaUpdateIOOFriday's In-Vac work
Liyuan recently did some of his pencil beam scatterometer measurements measuring not the
BRDF but instead the total integrated power radiated from each surface point
of some of the spare small optics (e.g. MMT, MC1, etc.).

The results are here on the iLIGO Wiki.

So some of our loss might just be part of the coating.
  61   Sun Nov 4 23:55:24 2007 ranaUpdateIOOFriday's In-Vac work
On Friday morning when closing up we noticed that we could not get the MC to flash any modes.
We tracked this down to a misalignment of MC3. Rob went in and noticed that the stops were
still touching. Even after backing those off the beam from MC3 was hitting the east edge of
the MC tube within 12" of MC3.

This implied a misalignment of MC of ~5 mrad which is quite
large. At the end our best guess is that either I didn't put the indicator blocks in the
right place or that the MC3 tower was not slid all the way back into place. Since there
is such a strong stickiness between the table and the base of the tower its easy to
imagine the tower was misplaced.

So we looked at the beam on MC2 and twisted the MC3 tower. This got the beam back onto the
MC2 cage and required ~1/3 if the MC3 bias range to get the beam onto the center. We used
a good technique of finding that accurately: put an IR card in front of MC2 and then look
in from the south viewport of the MC2 chamber to eyeball the spot relative to the OSEMs.

Hitting MC2 in the middle instantly got us multiple round trips of the beam so we decided
to close up. First thing Monday we will put on the MC1/MC3 access connector and then
pump down.


Its possible that the MC length has changed by ~1-2 mm. So we should remeasure the length
and see if we need to reset frequencies and rephase stuff.
  60   Sun Nov 4 23:22:50 2007 waldmanUpdateOMCOMC PZT and driver response functions
I wrote a big long elog and then my browser hung up, so you get a less detailed entry. I used Pinkesh's calibration of the PZT (0.9 V/nm) to calibrate the PDH error signal, then took the following data on the PZT and PZT driver response functions.:

  • FIgure 1: PZT dither path. Most of the features in this plot are understood: There is a 2kHz high pass filter in the PZT drive which is otherwise flat. The resonance features above 5 kHz are believed to be the tombstones. I don't understand the extra motion from 1-2 kHz.
  • Figure 2: PZT dither path zoom in. Since I want to dither the PZT to get an error signal, it helps to know where to dither. The ADC Anti-aliasing filter is a 3rd order butterworth at 10 kHz, so I looked for nice flat places below 10 KHz and settled on 8 kHz as relatively harmless.
  • Figure 3: PZT LSC path. This path has got a 1^2:10^2 de-whitening stage in the hardware which hasn't been digitally compensated for. You can see its effect between 10 and 40 Hz. The LSC path also has a 160 Hz low path which is visible causing a 1/f between 200 and 500 Hz. I have no idea what the 1 kHz resonant feature is, though I am inclined to point to the PDH loop since that is pretty close to the UGF and there is much gain peaking at that frequency.
Attachment 1: 071103DitherShape.png
071103DitherShape.png
Attachment 2: 071103DitherZoom.png
071103DitherZoom.png
Attachment 3: 071103LSCShape.png
071103LSCShape.png
Attachment 4: 071103DitherShape.pdf
071103DitherShape.pdf
Attachment 5: 071103DitherZoom.pdf
071103DitherZoom.pdf
Attachment 6: 071103LSCShape.pdf
071103LSCShape.pdf
Attachment 7: 071103LoopShape.pdf
071103LoopShape.pdf
  59   Sat Nov 3 16:20:43 2007 waldmanSummaryOMCA good day's work

I followed up yesterday's test of the PZT with a whole mess of characterizations of the PZT control and finished the day by locking the OMC with a PZT dither lock and a 600 Hz loop. I haven't analyzed any of the data yet, so its not calibrated in physical units and etc. etc. etc. Since a lot of the sweeps below are of a "drive the PZT, look at the PDH signal" nature, a proper analysis will require taking out the loop and calibrating the signals, which alas, I haven't done. Nonetheless, I include all the plots because they are pretty. The files included below are:

  • DitherLock_sweep: Sweep of the IN2/IN1 for the dither lock error point showing 600 Hz UGF
  • HiResPZTDither_sweep: Sweep of the PZT dither input compared to the PDH error signal. I restarted the front end before the sweep was finished accounting for the blip.
  • HiResPZTDither_sweep2: Finish of the PZT dither sweep


More will be posted later.
Attachment 1: 071103_DitherLock_sweep.png
071103_DitherLock_sweep.png
Attachment 2: 071103_DitherLock_sweep.pdf
071103_DitherLock_sweep.pdf
Attachment 3: 071103_HiResPZTDither_sweep.png
071103_HiResPZTDither_sweep.png
Attachment 4: 071103_HiResPZTDither_sweep.pdf
071103_HiResPZTDither_sweep.pdf
Attachment 5: 071103_HiResPZTDither_sweep2.png
071103_HiResPZTDither_sweep2.png
Attachment 6: 071103_HiResPZTDither_sweep2.pdf
071103_HiResPZTDither_sweep2.pdf
  58   Fri Nov 2 12:18:47 2007 waldmanSummaryOMCLocked OMC with DCPD
[Rich, Sam]

We locked the OMC and look at the signal on the DCPD. Plots included.
Attachment 1: 071102_OMC_LockedDCPD.gif
071102_OMC_LockedDCPD.gif
Attachment 2: 071102_OMC_LockedDCPD.pdf
071102_OMC_LockedDCPD.pdf
  57   Fri Nov 2 08:59:30 2007 steveBureaucracySAFETYthe laser is ON
The psl laser is back on !
  56   Thu Nov 1 20:03:00 2007 Andrey RodionovSummaryPhotosProcedure "Drop and Drag" in pictures
Attachment 1: DSC_0072.JPG
DSC_0072.JPG
Attachment 2: DSC_0083.JPG
DSC_0083.JPG
Attachment 3: DSC_0099.JPG
DSC_0099.JPG
Attachment 4: DSC_0100.JPG
DSC_0100.JPG
  55   Thu Nov 1 19:58:07 2007 Andrey RodionovBureaucracyPhotosSteve and Tobin's picture
Attachment 1: DSC_0023.JPG
DSC_0023.JPG
  54   Thu Nov 1 19:55:59 2007 Andrey RodionovBureaucracyPhotosAndrey, Tobin, Robert - photo
Attachment 1: DSC_0092.JPG
DSC_0092.JPG
  53   Thu Nov 1 19:55:03 2007 Andrey RodionovBureaucracyPhotosAndrey's photo
Attachment 1: DSC_0055.JPG
DSC_0055.JPG
  52   Thu Nov 1 19:54:22 2007 Andrey RodionovBureaucracyPhotosRana's photo
Attachment 1: DSC_0120.JPG
DSC_0120.JPG
  51   Thu Nov 1 19:53:34 2007 Andrey RodionovBureaucracyPhotosRobert's photo
Attachment 1: DSC_0068.JPG
DSC_0068.JPG
  50   Thu Nov 1 19:53:02 2007 Andrey RodionovBureaucracyPhotosTobin's picture
Attachment 1: DSC_0053.JPG
DSC_0053.JPG
  48   Thu Nov 1 16:51:33 2007 d40AoGGeneralD40
If you vant see D40 againn, you leave one plate goulash by N2 tank in morning.

Vit the good paprikash this time!!!
Attachment 1: PB010001.JPG
PB010001.JPG
  47   Thu Nov 1 16:42:48 2007 Andrey RodionovSummaryEnvironmentEnd of Daylight Saving Time this weekend
Useful information for everyone, as a friendly reminder:

According to the web-page

http://www.energy.ca.gov/daylightsaving.html,

this coming weekend there will be the end of Daylight Saving Time.

Clocks will be adjusted backward one hour.
  46   Thu Nov 1 16:34:47 2007 Andrey RodionovSummaryComputersLimitation on attachment size of E-LOG

I discovered yesterday when I was attaching photos that it is NOT possible to attach files whose size is 10Mb or more. Therefore, 10Mb or something very close to that value is the limit.
  45   Thu Nov 1 11:45:30 2007 tobinConfigurationIOOMode cleaner drag-wiping
Andrey, Bob, David, John Miller, Rana, Rob, Steve, Tobin

Yesterday we vented the vacuum enclosure and opened up the chamber containing MC1 & MC3 by removing the access connector between that chamber and the OMC chamber. Rana marked MC1's location with dogs and then slid the suspension horizontally to the table edge for easy drag-wiping access. The optic was thoroughly hosed-down with the dionizer, in part in an effort to remove dust from the cage and the top of the optic. Drag-wiping commenced with Rob squirting (using the 50 microliter syringe) and Tobin dragging (using half-sheets of Kodak lens tissue). We drag-wiped the optic many (~10) times, concentrating on the center but also chasing around various particles and a smudge on the periphery. There remains one tiny speck at about the 7:30 position, outside of the resonant spot area, that we could not dislodge with three wipes.

Today we drag-wiped MC3. First we slid MC1 back and then slid MC3 out to the edge of the table. We disconnected the OSEM cables in the process for accessibility, and MC1 is perched at an angle, resting on a dog. We did not blow MC3 with the deonizer, not wanting to blow particles from MC3 to the already-cleaned MC1. We drag-wiped MC3 only three times, all downward drags through the optic center, with Steve squirting and Tobin dragging. Some particles are still visible around the periphery, and there appears to be a small fiber lodged near the optic center on the reverse face.

Andrey and Steve have opened up MC2 in preparation for drag-wiping that optic after lunch.
  44   Thu Nov 1 09:17:27 2007 steveRoutineVACvent 64
Yesterday before vent I could not lock MC, therfore I could not measure the
transmitted power at MC2
The vent went well. We had lots of help.

We could not find the Nikon D40
PLEASE BORROW THINGS when taking them away
and bring them back promtly.

The laser was turned off for better visibility.

I see clean room frorks laying around here and there.
Please put them away so we do not carry excess particles into the chamber.
Attachment 1: vent64.jpg
vent64.jpg
  43   Thu Nov 1 01:28:04 2007 waldmanOtherOMCFirst digital lock of OMC
[Pinkesh, Sam]

We locked a fiber based NPRO to the suspended OMC tonight using the TPT digital control system. To control the laser frequency, we took the PZT AI output and ran it on a BNC cable down the hallway to the Thorlabs HV box. The Thorlabs is a singled ended unit so we connected the AI positive terminal only and grounded the BNC to the AI shield. We could get a -6 to 1.5 V throw in this method which fed into the 10 k resisotr + 9 V battery at the input of the HV box. The HV out ran to the NPRO PZT fast input.

We derived our error signal from a PDA255 in reflection with a 29.5 MHz PDH lock. The signal feeds into one of the unused Tip/Tilt AA channels and is passed to the PZT LSC drive through the TPT_PDH1 filter bank. In the PZT_LSC filter we put a single pole at 1 Hz which, together with the phase we mentioned the other night (180 degrees at 3 kHz) should allow a 1 kHz-ish loop. In practice, as shown below, we got a 650 Hz UGF with 45 degrees of phase margin and about 6 dB of gain margin.

The Lower figure shows the error point spectrum with 3 settings. REF0 in blue shows lots of gain peaking at 1.5 kHz-ish, just where its expected - the gain was -40. The REF1 has gain of -20 and shows no gain peaking. The current trace in red shows some gain peaking cuz the alignment is better but it also has included a 1^2:20^2 boost which totally crushes the low frequency noise. We should do a better loop sweep after getting the alignment right so we can see how much boost it will really take.

Just for fun, we are leaving it locked overnight and recording the PZT_LSC data for posterity.
Attachment 1: 071101_PZT_firstLoopSweep.pdf
071101_PZT_firstLoopSweep.pdf
Attachment 2: 071101_PZT_firstLoopSweep.gif
071101_PZT_firstLoopSweep.gif
Attachment 3: 071101_OMC_FirstLock_spectra.pdf
071101_OMC_FirstLock_spectra.pdf
Attachment 4: 071101_OMC_FirstLock_spectra.gif
071101_OMC_FirstLock_spectra.gif
  42   Wed Oct 31 23:55:17 2007 waldmanOtherOMCQPD tests
The 4 QPDs for the OMC have been installed in the 056 at the test setup. All 4 QPDs work and have medm screens located under C2TPT. The breadboard mounted QPDs are not very well centered so their signal is somewhat crappy. But all 4 QPDs definitely see plenty of light. I include light and dark spectra below. QPDs 1-2 are table-mounted and QPD 2 is labeled with a bit of blue tape. QDPs 3-4 are mounted on the OMC. QPD3 is the near field detector and QPD4 is the far field. In other words, QPD3 is closest to the input coupler and QPD4 is farthest.

Included below are some spectra of the QPDs with and without light. For QPDs 1 & 2, the light source is just room lights, while 3&4 have the laser in the nominal OMC configuration with a few mWs as source. The noise at 100 Hz is about 100 microvolts / rtHz. If I recall correctly, the QPDs have 5 kOhm transimpedance (right Rich?) so this is 20 nanoamps / rtHz of current noise at the QPD.
Attachment 1: QPD_SignalSpectrum.pdf
QPD_SignalSpectrum.pdf
Attachment 2: QPD_SignalSpectrum.gif
QPD_SignalSpectrum.gif
  41   Wed Oct 31 19:26:08 2007 Andrey RodionovRoutineGeneralPhotographs of "Mode-Cleaner Entrance"

Here are the pictures of "inside the chamber".
Attachment 1: MC-Pictures-1.pdf
MC-Pictures-1.pdf MC-Pictures-1.pdf MC-Pictures-1.pdf MC-Pictures-1.pdf
Attachment 2: MC-Pictures-2.pdf
MC-Pictures-2.pdf MC-Pictures-2.pdf MC-Pictures-2.pdf MC-Pictures-2.pdf
Attachment 3: MC-Pictures-3.pdf
MC-Pictures-3.pdf MC-Pictures-3.pdf MC-Pictures-3.pdf MC-Pictures-3.pdf
Attachment 4: MC-Pictures-4.pdf
MC-Pictures-4.pdf MC-Pictures-4.pdf MC-Pictures-4.pdf MC-Pictures-4.pdf
Attachment 5: MC-Pictures-5.pdf
MC-Pictures-5.pdf MC-Pictures-5.pdf MC-Pictures-5.pdf MC-Pictures-5.pdf
Attachment 6: MC-Pictures-6.pdf
MC-Pictures-6.pdf MC-Pictures-6.pdf MC-Pictures-6.pdf MC-Pictures-6.pdf
Attachment 7: MC-Pictures-7.pdf
MC-Pictures-7.pdf MC-Pictures-7.pdf MC-Pictures-7.pdf MC-Pictures-7.pdf
Attachment 8: MC-Pictures-8.pdf
MC-Pictures-8.pdf MC-Pictures-8.pdf MC-Pictures-8.pdf MC-Pictures-8.pdf
Attachment 9: MC-Pictures-9.pdf
MC-Pictures-9.pdf MC-Pictures-9.pdf
  40   Wed Oct 31 15:22:59 2007 robConfigurationIOOMode Cleaner transfer function
I measured the transfer function of the input mode cleaner using a PDA255 and the ISS. First I put the PD in front of the ISS out-of-loop monitor diode and used an SR785 to measure the swept sine transfer function from the Analog IN port of the ISS to the intensity at the PD. Then I moved the PD to detect the light leaking out from behind MC2, using ND filters to get the same DC voltage, and measured the same transfer function. Dividing these two transfer functions should take out the response of the ISS and the PD, and leave just the transfer function of the MC. A plot of the data, along with a single-pole fit, are attached.

The fit is pretty good for a single pole at 3.79 kHz. There's a little wiggle around 9kHz due to ISS weirdness (as Tobin has not been giving it the attention it requires), but this shouldn't affect this result too much. Using the known MC length of 27.0955m, and assuming that MC1 and MC3 have a power transmissivity of 2000ppm and MC2 is perfectly reflecting, the total round trip loss should be about 300ppm. The fitted finesse is 1460.
Attachment 1: MCtf.pdf
MCtf.pdf
  39   Wed Oct 31 15:02:59 2007 tobinRoutineIOOMode Cleaner Mode Tracking
I processed the heterodyned mode cleaner data yesterday, tracking the three 28 kHz modes corresponding to MC1, MC2, and MC3. Unfortuntately the effect of our MC power chopping is totally swamped by ambient temperature changes. Attached are two plots, one with the tracked mode frequencies, and the other containing dataviewer trends with the MC transmitted power and the room temperature. Additionally, the matlab scripts are attached in a zip file.
Attachment 1: mode-track.pdf
mode-track.pdf
Attachment 2: trends.pdf
trends.pdf
Attachment 3: mcmodetrack.zip
  38   Wed Oct 31 10:31:23 2007 Andrey RodionovRoutineVACVenting is in progress

We (Steve, David, Andrey) started venting the vacuum system at 9.50AM Wednesday morning.
  37   Wed Oct 31 09:45:28 2007 waldmanOtherOMCResolution to DAQland saga
[Jay, Sam]

We did a rough accounting for the linear delay this morning and it comes out more or less correct. The 10 kHz 3rd order butterworth AA/AI filter gives ~90 degrees of phase at 6 kHz, or 42 microseconds. Taken together, the two AA and AI filters are worth 80 microseconds. The 1.5 sample digital delay is worth 1.5/32768 = 45 microseconds. The remaining 160 - 125 = 35 microseconds is most likely taken up by the 64 kHz to 32 kHz decimation routine, assuming this isn't accounted for already in the 1.5 sample digital delay.

It remains to be seen whether this phase delay is good enough to lock the laser to the OMC cavity
  36   Wed Oct 31 08:38:35 2007 ranaProblem FixedIOOMC autolocker
The MC was having some trouble staying locked yesterday. I tracked this down to some steps in the last
half of the mcup script; not sure exactly which ones.

It was doing something that made the FAST of the PSL go to a rail too fast for the SLOW to fix.
So, I broke the script in half so that the autolocker only runs the first part. We'll need to
fix this before any CM locking can occur.

We also need someone to take a look at the FSS Autolocker; its ill.
  35   Wed Oct 31 08:34:35 2007 ranaOtherIOOloss measurements
In the end, we were unable to get a good scatter measurement just because we ran out of steam. The idea was to get a frame
grab image of MC2 but that involves getting an unsaturated image.

In the end we settle for the ringdowns, Rob's (so far unlogged) cavity pole measurement, and the MC transmission numbers. They
all point to ~100-150 ppm scatter loss per mirror. We'll see what happens after wiping.
  34   Wed Oct 31 08:33:54 2007 ranaProblem FixedSUSVent measurements
There was a power outage during the day yesterday; whoever was around should post something here about the
exact times. Andrey and David and Tobin got the computers back up - there were some hiccups which you can
read about in David's forthcoming elog entry.

We restarted a few of the locking scripts on op340m: FSSSlowServo, MCautolocker. Along with the updates
to the cold restart procedures we have to put an entry in there for op340m and a list of what scripts
to restart.

David tuned up the FSS Slow PID parameters a little; he and Andrey will log some entry about the proper
PID recipe very soon. We tested the new settings and the step response looks good.

We got the MC locking with no fuss. The 5.6 EQ in San Francisco tripped all of the watchdogs and I upped
the trip levels to keep them OK. We should hound Rob relentlessly to put the watchdog rampdown.pl into
the crontab for op340m.
  33   Tue Oct 30 20:15:24 2007 tobinOtherEnvironmentearthquake
Rana, Tobin

Largish (M5.6) earthquake in San Francisco sent our optics swinging.
  32   Tue Oct 30 19:32:13 2007 tobinProblem FixedComputersconlogger restarted
I noticed that the conlogger wasn't running. It looks like it hasn't been running since October 11th. I modified the restart_conlogger script to insist that it run on op340m instead of op440m, and then ran it on op340m.
  31   Tue Oct 30 16:55:40 2007 tobinRoutine Drag-wiping perfected
Steve, Tobin

Steve procured an assortment of syringes from the bio storeroom and we practiced drag-wiping the SOS in the flow bench. Using a 50 microliter Hamilton syringe to deliver 16 microliters of methanol seems perfect for drag-wiping the small optics. Drag-wiping in the downward direction seems to work very well, since we can squirt the optic directly in the center, and the (half) piece of kodak lens tissue fits easily between the bottom two earthquake stops.
  30   Tue Oct 30 13:58:07 2007 ajwConfigurationIOOMC Ringdowns
Here's a quick fit-by-eye to the latter part of the data from tek00000.xls.

The prediction (blue) is eqn 41 of
http://www.ligo.caltech.edu/docs/P/P000017-A.pdf

T1 = T2 = 0.002. Loss1 = Loss2 = 150 ppm.
MC3 assumed perfectly reflecting.
Velocity = 320 um/s (assumed constant), 2 usec into the ringdown.

OK, there's one little fudge factor in the prediction:
I multiplied D by 2.
Attachment 1: CavityRingdown.png
CavityRingdown.png
Attachment 2: CavityRingdown.m
% CavityRingdown.m
% Eqn 41 of 
% "Doppler-induced dynamics of fields in Fabry–Perot
% cavities with suspended mirrors", Malik Rakhmanov (2000).
% http://www.ligo.caltech.edu/docs/P/P000017-A.pdf

clear all

% read in ringdown timeseries:
at = importdata('tek00000.csv');
... 121 more lines ...
  29   Tue Oct 30 00:47:29 2007 ranaOtherIOOMC Ringdowns
I did a bunch of MC ringdown measurements using the PD that Rob set up. The idea is to put a fast PD (PDA255)
looking at the transmission through MC2 after focusing by a fast lens. The input to the MC is turned off fast
by flipping the sign of the FSS (Andri Gretarsson's technique).

With the laptop sitting on the MC can, its easy to repeat many ringdowns fast:
- Turn off the MC autolocker. Relock the MC with only the acquisition settings; no boosts
  and no RGs. This makes it re-acquire fast. Turn the MC-WFS gain down to 0.001 so that
  it keeps it slowly aligned but does not drift off when you lose lock.

- Use low-ish gain on the FSS. 10 dB lower than nominal is fine.

- Setup the o'scope (100 MHz BW or greater) to do single shot trigger on the MC2 trans.

- Flip FSS sign.

- Quickly flip sign back and waggle common gain to get FSS to stop oscillating. MC
  should relock in seconds.

Clearly one can scriptify this all just by hooking up the scope to the ethernet port.


Attached are a bunch of PNG of the ringdowns as well as a tarball with the actual data. A sugar
napoleon to whomever can explain the 7 us period of the wiggle before the vent!
Attachment 1: tek00000.png
tek00000.png
Attachment 2: tek00001.png
tek00001.png
Attachment 3: tek00004.png
tek00004.png
Attachment 4: MC2ringdown.tar.gz
  28   Mon Oct 29 23:25:42 2007 tobinSoftware InstallationCDSframes mounted
I mounted the frames directory on mafalda and linux3. It's intentionally not listed in the /etc/fstab so that an fb crash won't prevent the controls machines from booting. The command to mount the frames directory is:

mount fb40m:/frames/frames /frames
  27   Mon Oct 29 23:10:05 2007 waldmanConfigurationOMCLost in DAQspace
[Pinkesh, Sam]

In setting up a Digital based control of the hanging OMC, we naively connect the Anti-Imaging filter output to an Anti-Aliasing input. This led to no end of hell. For one thing, we found the 10 kHz 3rd order butterworth at 10 kHz, where it should be based on the install hardware. One wonders in passing whether we want a 10 kHz butter instead of a 15 kHz something else, but I leave that for a later discussion. Much more bothersome is a linear phase shift between output and input that looks like ~180 microseconds. It screams "What the hell am I!?" and none of us could scream back at it with an answer. I believe this will require the Wilson House Ghost Busters to fully remedy on the morrow.
Attachment 1: SS.pdf
SS.pdf
Attachment 2: SS.gif
SS.gif
  26   Mon Oct 29 12:20:15 2007 waldmanConfigurationOMCChanged OMS filters
I changed the OMS configuration so that some of the OMC-SUS LED channels go to a breakout box so that we can input the PDH error signal. After lunch, we will try to lock the cavity with a PDH error signal and digital filters. Then its on to dither locked stuff. Note that this LED business will have to be changed back some day. For now, it should be extremely visible because there are dangling cables and a hack job interface lying around.
  25   Mon Oct 29 11:07:22 2007 waldmanSoftware InstallationOMCSoftware install on OMS
[Alex, Sam]

We spent a little time this morning working on OMS and getting things restarted. A few changes were made. 1) We put openmotif on OMS so that the burtrb doesn't throw that crappy libXm any more. 2) We upgraded OMS to a 32 kHz sampling rate from 2 kHz. All the filters will have to be changed. We also added a PDH filter path to maybe feedback PDH signals cuz that will be cool. Maybe someday I will write up the very cool channel adding procedure.
  24   Mon Oct 29 09:46:50 2007 steveRoutineVACvac & pem trend
Pumpdown 64 pumped by maglev for 125 days
pd64-m-d125

Rob, can you tell me, when did the fire start on this plot?
Attachment 1: pd64md125.jpg
pd64md125.jpg
  23   Mon Oct 29 09:16:31 2007 steveRoutineVACthe rga is back
We had no filament current since last power glitch of Oct. 8, 2007
First I thought that the filament was lost, but it was only bad contact.
The rga head pins were oxidized. Rga was turned back on last Friday.
It's temp is 55.3C normal
Attachment 1: rgaisback.jpg
rgaisback.jpg
  22   Sun Oct 28 03:03:42 2007 ranaConfigurationIOOThree Way Excitement
We've been trying to measure the MC mirror internal mode frequencies so that we can measure
their absorption before and after drag wiping.


It looked nearly impossible to see these modes as driven by their thermal excitation level;
we're looking at the "MC_F" or 'servo' output directly on the MC servo board.

Today, I set up a band limited noise drive into the 'Fast POS' inputs of the 3 MC coil
driver boards (turns out you can do this with either the old HP or the SR785).

Frequencies:
MC1     28.21625 kHz
MC2     28.036   kHz
MC3     28.21637 kHz

I don't really have this kind of absolute accuracy. These are just numbers read off of the SR785.

The other side of the setup is that the same "MC_F" signal is going into the SR830 Lock-In which
is set to 'lock-in' at 27.8 kHz. The resulting demodulated 'R" signal (magnitude) is going into
our MC_AO channel (110B ADC).

As you can see from the above table, MC1 and MC3 are astonishingly and annoyingly very close in
frequency. I identified mirrors with peaks by driving one at a time and measuring on the spectrum
analyzer. I repeated it several times to make sure I wasn't fooling myself; it seems like they
are really very close
but distinct peaks. I really wish we had chipped one of these mirrors
before installing them.



Because of the closeness of these drumhead modes, we will have to measure the absorption by making long
measurements of this channel.
  21   Sat Oct 27 19:00:44 2007 waldmanConfigurationOMCHanging, locked OMC with REFL extracted.
I got the OMC locked to the fiber output today. It was much more difficult than I expected and I spent about 30 minutes or so flailing before stopping to think. The basic problem is that the initial alignment is a search in 4-dimensional space and there is naturally only one signal, the reflected DC level, to guide the alignment. I tried to eyeball the alignment using the IR card and "centering" the beams on mirrors, but I couldn't get close enough to get any light through. I also tried to put a camera on the high reflector transmission, but with 1.5 mW incident on the cavity, there is only 1.5 microwatts leaking through in the best case scenario, and much, much less during alignment.

I resolved the problem by placing a high reflector on a 3.5 inch tall fixed mount and picking off the OMC transmitted beam before it reaches the DC diodes. I took the pickoff beam to a camera. The alignment still sucked because even though the beam cleanly transmitted the output coupler, it wasn't anywhere close to getting through the OTAS. To resolve this problem, I visually looked through the back of M2 at M1 and used the IR card to align the beam to the centers of each mirror. That was close enough to get me fringes and align the camera. With the camera aligned, the rest was very easy.

I restored the PDH setup we know and love from the construction days and locked the laser to the OMC with no difficulty. The laser is in Rana's lab so I send the +/- 10V control signal from the SR560 down a cable to 058E where it goes into the Battery+resistor box, the Throlabs HV amplifier, and finally the FAST channel of the NPRO. BTW, a simple experiment sows that about 35 +/- 3 V are required to get an FSR out of the NPRO, hence the Thorlabs HV. The EOM, mixer, splitter, etc is on the edge of the table.

With this specific OMC alignment, ie. the particular sitting on EQ stops, it looks like all of the ghost beams have a good chance of coming clear. I can fit a 2 inch optic in a fixed mount in between the end of the breadboard and the leg of the support structure. A picture might or might not be included someday. One of the ghost beams craters directly into the EQ stop vertical member. The other ghost barely misses M2 on its way down the length of the board. In its current configuration, the many REFL beam misses the leg by about 1.5 inches.
  20   Fri Oct 26 21:48:40 2007 waldmanConfigurationOMCFiber to 056
I set up a 700 mW NPRO in Rana's lab and launched it onto a 50m fiber. I got a few mW onto the fiber, enough to see with a card before disabling the laser. The fiber now runs along the hallway and terminates in rm 056. Its taped down everywhere someone might trip on it, but don't go out of your way to trip on it or pull on it because you are curious. Tomorrow I will co-run a BNC cable and attenuate the NPRO output so it can only send a few mW and so be laser safe. Then we can try to develop a procedure to align the beam to a suspended OMC and lock our suspended cavity goodness.

Notes to self: items needed from the 40m
  • ND10 and ND20 neutral density filter
  • EOM and mount set for 4 inch beam height
  • Post for fiber launch to get to 4 inch
  • Mode matching lens at 4in
  • 3x steering mirror at 4in
  • RF photodiode at 4in
  • Post for camera to 4in
  • Light sheild for camera
  • Long BNC cable
Some of these exist at 056 already
  19   Fri Oct 26 17:34:43 2007 waldmanOtherOMCOMC + earthquake stops

[Chub, chris, Pinkesh, Sam]

Last night we hugn the OMC for the first time and came up with a bunch of pictures and some problems. Today we address some of the problems and, of course, make new problems. We replaced the flat slotted disks with the fitted slotted disks that are made to fit into the counterbore of the breadboard. This changed the balance slightly and required a more symmetric distribution of mass. It probably did not change the total mass very much. We did find that the amount of cable hanging down strongly affected the breadboard balance and may also have contributed to the changing balance.

We also attached earthquake stops and ran into a few problems:

  • The bottom plate of the EQ stops is too thick so that it bumps into the tombstones
  • The vertical member on the "waist" EQ stops is too close to the breadboard, possibly interfering with the REFL beam
  • The "waist" EQ stops are made from a thin plate that doesn't have enough thickness to mount helicoils in
  • Helicoil weren't loaded in the correct bottom EQ stops
  • The DCPD cable loops over the end EQ stop looking nasty but not actually making contact

However, with a little bit of jimmying, the EQ stops are arrayed at all points within a few mm of the breadboard. Meanwhile, Chub has cabled up all the satellite modules and DCPD modules and Pinkesh is working on getting data into the digital system so we can start playing games. Tonight, I intend to mount a laser in Rana's lab and fiber couple a beam into the 056 room so we can start testing the suspended OMC.
  18   Fri Oct 26 16:19:29 2007 Tobin FrickeRoutineIOOMC resonances
We would like to measure the absorption of the mode cleaner optics. The plan is to repeat <a href="http://ilog.ligo-wa.caltech.edu:7285/mLIGO/Cleaning_the_Mode_Cleaner">Valera's experiment</a> in which we track the MC's thermal resonances to infer their power absorption. Last night Rana and I hooked up a lock-in amplifier to heterodyne the MC servo signal by 28 kHz and piped the output into an ADC using the MC_AO channel. We did not find any resonances.

Valera recommends we drive the POS of the three MC optics with bandlimited noise to excite the resonances.
  17   Fri Oct 26 09:10:17 2007 steveRoutinePEMPEM &PSL trend
The fires are out, lab particle counts are up.
Psl HEPAs are at 100% and mobel HEPAs are just turned on
20 days plot and 5 hrs plot below
Attachment 1: counts&psl.jpg
counts&psl.jpg
Attachment 2: 5dcounts.jpg
5dcounts.jpg
  16   Thu Oct 25 23:35:36 2007 waldmanOtherOMCHang the OMC!
[Pinkesh, Sam]

We tried, convicted and hung the OMC today. The OMC was found guilty of being overweight, and unsymmetrically balanced. The unsymmetry was kind of expected and was corrected with a hefty stack of counterweights positioned over the counterweighting holes. The stacks will be measured at some future date and correctly sized objects machined. The overweightness showed up when the level hanging breadboard was about 5 mm low. This showed up in the board height above the table as well as the OSEM flag positions within their holes. The problem was remedied with a liposuction of the intermediate mass. We removed both small vertical cylinder weights that Chris added, and then we removed the heavy steel transverse weight that can be used to adjust the tip around the long axis (I forgot what its called).

The top of the breadboard ended up about 154 mm off the table. The breadboard is 39 mm thick, and the optics are centered (30 - 12.7) = 17.3 mm below the surface for a as hanging beams height of 154 -39 - 17.3 = 97.7 mm or about an 0.150 inches lower than we were aiming for. Can I get a refund?

We screwed up in multiple ways:
  • The slotted disks that capture the wires do not have the alignment bore used to center the wire in the hole
  • We didn't correctly route the far field QPD cable so it runs funny
  • We didn't have a tool which could be used to get two of the DCPD preamp box mounting screws (which are M3's chub!)
  • We don't have the cable clamps to tie off the electrical cables to the intermediate mass
  • We don't have any of the cabling from the OMC-SUS top to the rack so we can't test anything
  • We haven't uploaded pretty pictures for all to see

We left the OMC partially suspended by the OMC-SUS and partly resting on the installation lab jacks which are currently acting as EQ stops. After we fix the cabling we will more permanently hang it. PS, It looks like the REFL beam extraction will be tricky so we need to get on that....
Attachment 1: IMG_1483.jpg
IMG_1483.jpg
Attachment 2: IMG_1481.jpg
IMG_1481.jpg
  15   Thu Oct 25 22:02:58 2007 robRoutinePSLHEPAs maxed
In light of the SoCal fires, I turned the PSL HEPAs up to 100%.
  14   Thu Oct 25 17:52:45 2007 waldmanOtherOMCOMCs with QPDs
[Rich, Chub, Pinkesh, Sam]

Yesterday we got the QPD, OTAS, and PZT cabling harness integrated with the OMC. We found a few things out, not all of them good. The QPDs went on no problem and could be fairly well aligned by hand. We "aligned" them by looking at all four channels of the QPD on the scope and seeing that there is signal. Since the beam is omega = 0.5 mm, this is a reasonable adjustment. We then connected the OTAS connector to the OTAS and found that the heater on the OTAS was bonded on about 30 degrees rotated from its intended position. This rotated the connector into the beam and caused a visible amount of scattering. This wasn't really a disaster until I removed the connector from the heater and broke the heater off of the aluminum parts of the OTAS. Two steps backwards, one step forward. After the OMC, OMC-SUS integration test we will re-bond the heater to the aluminum using VacSeal. In the meantime, the OMC has been moved to Bridge 056 for integration with the OMC-SUS. More on that as we make progress.
ELOG V3.1.3-